(-2,1); perpendicular to 3x+y=1
Answer by ikleyn(52778)

.

The line perpendicular to 3x+y = 1 is

    x - 3y = c,      (1)  where c is an arbitrary constant.


To specify the perpendicular line through the point (-2,1), put in equation (1)
x= -2,  y= 1.  You will get then

    c = -2 -3*1 = -2 - 3 = -5.


Thus the equation you are seeking for is  x - 3y = -5.    ANSWER

Solved.

-----------------

For more details and other forms of reasoning,  see the lesson
    - Equation for a straight line perpendicular to a given line and passing through a given point
in this site.